40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 10 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  2334   Sun May 5 20:52:16 2019 ranaNoise Budget2micronLasersFrequency noise measurement of 2 micron source

Looks good! A few comments:

  1. The LPF should be a 1.9 MHz LP from Mini-circuits
  2. The SR560 should be set with a 1 MHz low pass only
  3. No Rubidium needed. We just want to use two IFR/Marconi synthesizers so that they can share the same time standard.
  4. We will need a RF Amp to drive the AOM hard enough - check the manual and see what level is required to get the max diffraction.
  2333   Sun May 5 15:17:33 2019 AnjaliNoise Budget2micronLasersFrequency noise measurement of 2 micron source

We are going to start the heterodyne configuration for the frequency noise measurement of 2 micron laser source. The schematic of the same is shown in attachment #1. The initial plan was to separate the photodetector output to in phase , quadrature component and then extract the frequency noise information from the same. The new plan suggested is to use a phase locked loop for the demodulation part. We can use IFR Marconi 2023A as a VCO and SR560 as the amplifier. The RF generator for the AOM and the VCO need to be locked through Rubidium frequency standard (FS725). One of the issues is that the band width of the photodetector that we have is 14 MHz and the AOM is at 80 MHz. The response of photodetector at 80 MHz is about 20 % of that at DC.I have modified the noise budget after discussing with Anchal. Attachment #2 shows the noise budget. The Marconi data is taken from Anchal. The data plotted in attachment #2 corresponds to Marconi #539 at 48 MHz and actuation slope of 500 Hz. We may have to find the actual numbers based on our settings for the experiment.  From the data sheet, the input noise of SR 560 is 4nV/rtHz. The frequency mixer (Minicircuit ZAD-1-1+) has a conversion loss of 4.83 dB and the thermal noise of it is not given in the data sheet and thus it is not included in the noise budget. The length of the fiber considered for MZI is 10 m.

 

Attachment 1: schmeatic_heterodyne.png
schmeatic_heterodyne.png
Attachment 2: Noise_budget_modified.pdf
Noise_budget_modified.pdf
  2332   Wed May 1 14:34:53 2019 awadeDailyProgressWOPOOutput chain and demodulation of squeezing signal

The detectors have been verified to have a low intercept current, that will give us good clearance of our shot noise to dark noise (at least as measured in the <100 kHz band).  The question is if we can make a ~ 1 Mhz measurement of squeezing without extra noise in that readout chain dominating over our expected 1 dB squeezing.

The plan is to use the Zurich box (model HF2LI) to mix down quantum enhancement from 1 MHz in our homodyne detector.  This will be well above most of the classical noise of the laser and other acoustic pickup from the environment.  

The Zurich box has a input referred noise of 5 nV/rtHz above 10 kHz and operating at at the higher end of its sampling rate. It has a couple of nice features that would make it well suited to the task.  We can use the lock in amplifier instrument to digitally mix down signal from 1 MHz and output one or both of the i/q quadratures to its digital output ports.  From this instrument we can also set the bandwidth of the 

Output noise of Zurich is 25 nV/rtHz, so I'll need some gain factor to boost signal if readout is done from the external ports. Relative gain of the two paths can be changed manually by typing values into boxes, but might be tought to do a direct ballenced subtraction within the box live if the real time drift is large.

 

Laser diode modulation port on Diabolo has a signal bandwidth of 5 kHz, so can't really be used to generate a reference signal.  

  2331   Tue Apr 30 19:52:14 2019 awadeHowtoWOPOConfiguring Zurich box server

As I couldn't find this on the 40m elog/wiki I'm putting notes here on how to configure a fresh webserver + python server for the Zurich box (model HF2LI) that runs an autobooting service job. 

The Zurich box (model HF2LI) has no front panel interface.  All controls are done from a USB interface to a host computer (running windows or Ubuntu/Debian).  This computer then hosts a web-server that runs an interface in a browser (through http) or responds to requests from python/matlab/labview/C API over the network.  

Installing data and webserver 

Set up on a Linux box is fast and easy from a fresh install.  A thin low powered computer is ideal. Start with a fresh install of Debian, use the latest stable release. Then download the latest stable release for the HF2LI, HF2IS, HF2PLL instruments from the Zurich downloads page.  You’ll probably want the 64bit linux. You can untar this file with

> tar xzvf LabOneLinux<arch>-<release>-<revision>.tar.gz

(Insert your filename details above)

Then cd into the directory it created and run

> sudo bash install.sh

It will ask a bunch of questions, just hit enter and y all the way through. The install should immediately boot the HF2 data server at the end.  You can check the data server is running with

> ziService status

Now plug in your Zurich box using the USB cable. Note: there are ethernet ports on the box, but these are NOT for network ethernet.  If the box is powered up the data server should find the box.  

Running the web server

In another terminal you’ll want to to start a Webserver so you can view the box status, settings and live data.  Go to another terminal and run

> ziWebServer

Your Zurich device should now be exposed on port 8006 of that machine.  You are read to go.

Now on that computer you can got to http://localhost:8006, alternatively on any other computer on the network you can go to the host machine’s IP address and see your Zurich interface.  I.e. if the computer IP is 192.168.1.121 then go to http://192.168.1.128:8006.  

Remote access

If you wanted to access the Zurich interface remotely (outside network) you can do some SSH forwarding if you have a gateway machine to outside the lab.  The standard command for this is 

> ssh -nNT -L <clientport>:localhost:<hostport> <controls>@<Host-IP-Address>

where controls is the username, Host-IP-Address is the IP address of the machine to be forwarded from, clientport is the port on your machine (the one you are ssh'ing from) and hostport is the port on the remote machine (for Zurich this is 8006).  Then when you go to http://localhost:<clientport> and you'll see Zurich interface.

Python API

Since the release 2.18 (Nov 2018) the python access tools are now just in Pypi.  You can install them with pip:

> pip install --upgrade zhinst

This should now allow you to use python API to access your Zurich box.

 

Setting up service to keep host alive from when box is booted

You’ll want to set up a service on your Debian box with systemd.  Make a unit

> cd /etc/systemd/system

> sudo vim ziServer.service

You must use sudo to be able to save the file here.

Paste the following in, this assumes that your username and its group is called controls on your machine:

[Unit]
Description=Zurich Instruments HF2LI data server
After=syslog.target network.target

[Service]
User=controls
Group=controls
WorkingDirectory=/home/controls/services/
ExecStart=/bin/bash -c "exec ziServer"
Restart=no
RestartSec=30

[Install]
WantedBy=multi-user.target


Tip: in vim you can type “:set paste” to get exact paste without auto indent etc.

Save the file and exit.  You'll also need to create the working director, or change the above to your desired directory

> mkdir /home/controls/services/

Finally, tell systemd to look for new things:

> systemctl daemon-reload

and enable the service,

> systemctl enable ziServer.service

and manually start the service,

> systemctl start myservice.service

It should be running, try to see if this is so:

> ps | grep ziServer

It should run on reboot of the machine and restart every 30 seconds if it crashes for whatever reason. Try rebooting the host machine to check it is still running

Configuring web-server service

Follow the above steps to also auto load the webserver with a unit file named /etc/systemd/system/ziWebServer.service:

[Unit]
Description=Zurich Instruments HF2LI web server
After=syslog.target network.target

[Service]
User=controls
Group=controls
WorkingDirectory=/home/controls/services/
ExecStart=/bin/bash -c "exec ziWebServer"
Restart=no
RestartSec=30

[Install]
WantedBy=multi-user.target

 Now you have a local network ready server for accessing the Zurich box through a web browser.

  2330   Wed Apr 24 19:30:55 2019 ranaLab InfrastructureHVAChowto get temperature trends

We need to get the temperature data from the sensors that Aidan setup. I have got it by downloading a sketchy HOBO app onto my phone. In order for this to work you have to go stand next to the sensors and download via bluetooth. There is no "pairing" required - the app seems to recognize the sensors nearly immediately.

and please, Aidan and everyone else, please do not use double-sticky tape or glue or bubble gum to attach sensors in a clean optical setup. We're going to have to make a mess to remove these.

I've just downloaded more data (attached). The lab seemed sort of cold and the south HEPA output was blowing colder air than the north HEPA output.

I've changed the temperature from 70F to 73F at 7:45 PM today. Lets see how it goes.

Attachment 1: QIL_temperature_190424.tgz
  2329   Wed Apr 24 14:19:55 2019 awadeLab InfrastructureHVACHVAC QIL: two actuators tied to one sensor now

RXA: I talked with Modesto this morning. He had gotten the information at some point that our only problem was that we had 2 thermostats, and had not been told about the weird temperature fluctuations and the hot air. So I talked with him this morning so that he knows that we're going for stability in addition to using 1 sensor.

AWADE (Additional Info Wed Apr 24 19:42:32 2019 ): Modesto said that he set the temperature to 70 F.  

Modesto returned today to carry out the fix of the AC.  The south side hot actuator on the south vent was tied to the west wall thermostat.  So current status is that both heat/cooling actuators are tied to the one thermostat on the west wall.  Work was completed at 2 pm.

Modesto said he would be back tomorrow to check the system.

Quote:

Someone from Facilities named Modesto came out to the QIL to see the problem with the HVAC. He measured 80F coming out of the hot vent (when its thermostat was set to 68F) and 65F coming out of the cold vent. He was able to drop the hot vent to about 62F by turning the set point down to 55F (so the controller is at least working).

I asked that he set it up that both units are controlled by one temperature sensor. So he’s going to disconnect the hot vent controller from it’s thermostat and tie it into the thermostat for the cold side. This seems reasonable but could be potentially problematic if the HVAC units have significantly different responses.

Anyway, the current plan is for him to come in on Friday to do this work

 

  2328   Wed Apr 24 09:01:49 2019 AnjaliUpdate2micronLasersIntensity noise stabilization of 2 micron source

Attachment #1 shows the schematic of the experimental setup for amplitude stabilization using AOM. The proposed idea is as follows

  • The output power from the  Zeroth order port of the AOM depends on the RF power supplied to the AOM and the  RF power from the RF driver is dependent on the input voltage to the driver. 
  • When amplitude of the laser chnages to due to amplitude noise, the output from photodiode will also changes. The output of one of the photodiodes is connected through a gain element (G) to the RF driver of AOM. So, when PD output changes , the gain of the gain element should be adjusted such that the input voltage to the AOM driver changes and thus the RF power which affect the output power from the AOM. 
  • The gain element could be SR560 (I still have to understand more about it) and the photodiodes that we have are thorlabs DET10D detector which are DC coupled. We also need additional TIA stage after the PD. There were two of them built by the summer student. We tested both and one of them only works fine now.
Quote:

OK...but how will the amplitude stabilization be done? How about a diagram showing the feedback loop and electronics?

  1. what is the feedback circuit?
  2. Can it drive the low impedance of an AOM driver?
  3. What is the photodiode?
  4. It has to be an AC coupled feedback (DC coupled is always terrible).
  5. What is the requirement on the intensity noise as a function of frequency?
Quote:

[Aidan, Anjali]

  • Attachment #1 shows the cleaned up setup for  2 micron experiment

 

Attachment 1: Intensity_stabilisation.png
Intensity_stabilisation.png
  2327   Mon Apr 22 19:35:16 2019 anchal and awadeSummaryWOPOShot Noise Intercept Current measurement of Photodiodes

Andrew measured the output voltage noise of photodiodes through an AC coupled pre-amplifier (G=200, 0.7 nV/rtHz) while varying the intensity of white light falling on it. The setup uses a regular 2xAA flashlight (torch) with an incandescent bulb to illuminate the photodiode. The flashlight focus is adjusted to give a non-expanding beam and a 25.4 mm (1") lens is then used to focus the collimated beam down. The total power of light was controlled by trimming it with an iris.


Data Analysis:

Below is the fitting of data to estimate the shotnoise intercept current. Here the fitted function is the quadrature sum of input referred shot noise and dark noise modeled as an equivalent shot noise quanitiy i_snint (called intercept current):

i_\textrm{in} = \sqrt{2e^-(i_\textrm{pc}+i_\textrm{snint})}

I have included zip file with data and notebook in it as well. With measured shotnoise intercept current of 48 ± 7 uA and 31 ± 7 uA, we are good to go to use these for measuring shot noise with light power in the order of 5.0 mW falling on the photodiodes. With 5.0 mW (~3.9 mA) on each PD, this gives us a clearance of 10dB from the dark noise floor on average for each detector.  Plots with fits are attached below.

Attachment 2: WOPO_BHD_Photodiodes_SNInterceptAnalysis.pdf
WOPO_BHD_Photodiodes_SNInterceptAnalysis.pdf WOPO_BHD_Photodiodes_SNInterceptAnalysis.pdf
Attachment 3: 20190419_MeasuredInterceptCurve.zip
  2326   Thu Apr 11 20:50:17 2019 ranaHowtoGeneralhow to learn quickly

https://maglab.caltech.edu/human-magnetic-reception-laboratory/

  2325   Wed Apr 10 18:58:48 2019 ranaUpdate2micronLasersIntensity noise stabilization of 2 micron source

OK...but how will the amplitude stabilization be done? How about a diagram showing the feedback loop and electronics?

  1. what is the feedback circuit?
  2. Can it drive the low impedance of an AOM driver?
  3. What is the photodiode?
  4. It has to be an AC coupled feedback (DC coupled is always terrible).
  5. What is the requirement on the intensity noise as a function of frequency?
Quote:

[Aidan, Anjali]

  • Attachment #1 shows the cleaned up setup for  2 micron experiment
  2324   Wed Apr 10 13:10:45 2019 awadeDailyProgressWOPORebuild of WOPO homodyne detectors using AD829

[awade]

This post details the rebuild I made of the transimpedance amplifiers (TIA) for the homodyne detectors.  My previous iteration of this design failed to take into account the output swing limitations and current draw limitations at the output of the op amp.  It turns out that the OP27 was not the best choice here.  On the face of it its GBP and input referred noise were fine for a ~ 1 MHz detection.  However, operating the chip with 5V output voltage (to get 10 dB shot noise clearance from thermal noise) and with output impedance on order of 150 Ω meant that the frequency response was not as expected (see ATF:2316). 

A better choice is something like AD829 which is designed for driving signals directly into 50Ω terminated outputs.  The updated design is illustrated below.

Here the transimpedance gain has been lowered to 2 kΩ to lower the relative input referred noise contributed by the AD829's slightly higher current noise.  Here the optical power can be increased to give the same effective output signal.  In this case increase optical power and lowering gain keeps the thermal noise clearance below shot noise constant but improves the clearance of the AD829 input current noise.

 

Scaling of thermal noise relative to shot noise given DC output voltage

For reference the thermal noise of the TIA is given by the Johnson noise

\delta n_\mathrm{thR} = \sqrt{\frac{4k_BT}{R_\textrm{fb}}}

where k_B is the  Boltzmann's constant, T is the temperature (300K) and Rfb is the feedback resistor value.  Also here the shot noise is

\delta n_\mathrm{SN} = \sqrt{2 e^- \mathcal{R}P_\textrm{PD}}

Where e is the charge on an electron (1.602e-19 C), R is the photodiode responsivity (~0.8 A/W), and P_PD is the power on the photodetectors (0.5-5 mW).  Keeping in mind that there is a maximum voltage swing (V_max) on the output of the op amp, this comes from its current driving capability into a given load, then the feedback resistance (TIA gain) is limited to 

R_\textrm{fb} = \frac{V_\textrm{Max}}{\mathcal{R} P_\textrm{PD}}

The noise clearance, given a maximum output voltage, is then given by the ratio

NR = \frac{\delta n_\mathrm{SN}}{\delta n_\mathrm{thR}} = \sqrt{\frac{e^-V_\textrm{Max}}{2kT}}

Power on the photodiode and responsivity cancel out for the noise clearance.  Assuming we are not optical power limited then the thermal noise ratio (clearance below shot noise) is wholly determined by V_Max and the physical constants e, k_B and T.  For room temperature (300 K) this reduces to the rule of thumb 

NR \sim 4.4 \sqrt{V_\textrm{Max}}

Keeping the V_max to within 3 V means that we will get a thermal noise clearance factor of 7.6 below shot noise (8.8 dB). V_max of 5.2 volts will give 10 dB clearance. Ok I guess. 

Choosing AD829

Koji suggested the AD829 as a replacement to OP27.  AD829 has applications in driving 50 Ω/75 Ω loads in video applications and has some nice noise characteristics.  The bottom line is that it has 1.7 nV/√Hz input voltage noise, 1.5 pA/√Hz input current noise, can do a ±3 V voltage swing into 150 Ω load (DC coupled), and is fast (600 MHz uncompensated, with 230 V/µs slew).  There are some quirks.  AD829 seems have have a weird 80 MHz feature that causes oscillations if not compensated properly.  I did a bit of modeling in LISO and then just decided to build it once I found that ~2 kΩ gain was about right for ensuring that the dark noise wasn't dominated by the op amp current noise.

Because I was building on proto-board I wanted to ensure that there was as little parasitic capacitance as possible.  I built the whole amplifier + PD onto a single SOIC-14 to 0.1 pitch adaptor (Adafruit SMT Breakout PCB for SOIC-14, Part No. 1210), pictured below.  The small footprint and short trace lengths between critical pins means that there is a lower chance of parasitic impedance causing instability or impact on bandwidth.  

Top view of Unit B TIA with PD mount soldered directly to inverting pin and positive bias
TIA at earlier stage of construction for both unit A and unit B
shows feedback resistor and capacitor soldered on reverse side (these passive
components are stacked on top of each other).

The feedback resistor (2 kΩ) and feedback compensating capacitor (15 pF ceramic + 3 pF nominal pin parasitic) were soldered directly to the reverse side where the TSSOP-14 pins were connected through vias to the top side.  This minimized the path length of this electrical signal.  I scratched off pads that were unused in case they shorted or were a source of crosstalk.  The photodiode mount pin was soldered directly to the edge of the board to minimize distance.  The power bypassing capacitors were SMT 100 nF ceramics (5%) that were soldered directly to the SMT breakout with the shortest path to ground on that board (see pictured).  The other photodiode pin was soldered to +5 V (-5V) supply pin in unit A (unit B).  The opposite biasing makes subtracting the signals using a summing circuit easier but may affect the overall TF response. There is no power regulation, I'm planning to use batteries to power these circuits.

In my initial tests I had no compensating capacitor attached to pin 5.  I inserted a 40 pF ceramic capacitor in place of the photodiode and looked at the AD829 output pin with a high impedance probe on an oscilloscope.  I immediately saw 80 MHz oscillations.  I don't really need a super high bandwidth so I went strait to the maximum recommended choice for pin5 compensating capacitor of 68 pF.  This makes AD829 unity gain stable with 66 MHz bandwidth (slew 16V/µs) but is more than enough for my needs.  This killed the high frequency RF ringing junk. Maybe less capacitance here would have worked, but this performance is enough for my needs and gives some certainty about the op amps stability (ignoring input capacitance compensation).

At the output I added 100 Ω of series resistance to limit the loading on the op amp and current draw when 50 Ω terminated.  With 50 Ω terminating impedance this makes a 150 Ω to ground.  Providing the output DC swing is kept within 3V the op amp should behave as expected.

The whole thing fits together very nicely on a single SMT breakout board. Making it compact will hopefully avoid any issues with stray capacitance messing up the performance.  I have then mounted this on a larger proto board for ease of installing in the experiment. Paths to the output SMA output are thin wire and not routed through the underlying board. Only power and ground is routed through the board, all other pins to the op amp disconnected. 

 

Transfer function PD

I measured the signal transfer function of the above TIA units using the Jenne rig at the 40 m.  It took me a number of tries: I made the mistake of loading the DC port of the NF 1611 detector with 50 ohms (which resulted in too much current draw and affected the LF response of the witness detector); there were also some issues with the polarization/positioning onto the beam splitter there that made it not 50:50 (it was 69:33 ), I moved the fiber launch a bit to get the splitting right and retook measurements after that; also I initially had the current set too high on the laser current driver which I think was saturating its response.  I'll skip over the details and just present the final measurement. I've included a zip in the attachments with the data that logs some more of the details of my failed attempts.

I used the same calibration as used in PSL:2247.  Here the laser diode current was set to 25.0 mA, which is 1.51 mW of 1064 nm, and the AC excitation into the laser was set to -21 dBm (19.9 mV rms).  Power on the witness detector was measured to be 0.79 mW and that on the PD under test was 0.75 mW (as measured with a power meter).  Units of measurement were linear magnitude and phase in degrees. DC voltages were measured with 1 MΩ impedance.

Unit A: using 25 mA current into laser and -21 dBm sweep I get DC on PD of 1.20 V (into 1 MΩ) and a DC voltage of NF1611 detector of -2.13 V (into 1 MΩ).

Unit B: using 25 mA current into laser and -21 dBm sweep I get DC on PD of 1.07 V (into 1 MΩ) and a DC voltage of NF1611 detector of -2.12 V (into 1 MΩ). 

The calibrated TIA current to voltage gain is plotted below:

A notebook detailing all the attempts is attached below.

 

 

Quote:

This set of measurments turned out to be bad. The OP27 (±15V supplies) with 50 Ω series at the output + 50 Ω load impedance seemed to be saturated.

Attachment 1: BasicSchematic_TranImpPD_AD829.pdf
BasicSchematic_TranImpPD_AD829.pdf
Attachment 4: WOPO_HDPD_AD829_2kohmTranImpGain_UnitAandBandLISO.pdf
WOPO_HDPD_AD829_2kohmTranImpGain_UnitAandBandLISO.pdf
Attachment 5: 20190405_TFHDPDs_AD829Design_LargeBoard.zip
  2323   Wed Apr 10 12:17:34 2019 AnjaliUpdate2micronLasersIntensity noise stabilization of 2 micron source

[Aidan, Anjali]

  • Attachment #1 shows the cleaned up setup for  2 micron experiment
  • We also have received the AOM . The initial test is to do the intensity stabilisation of 2 micron laser diode using AOM
  • Today, we started with the laser and AOM characterisation
  • Attachment #2 shows the schematic of the experimental setup for the laser characterisation. Laser output is connected to a Faray isolator and the isolator output is measured using Thorlabs thermal power meter.
  • The current limit on the current controller is set to 95 mA (maximum allowed current to the laser diode is 120 mA) and thermistor is set to 10 kohm (maximum value of thermistor resistance is 10 kohm)
  • Attachment #3 shows the output power as a function of input current to the laser diode.
  • From the data sheet of the laser diode, the threshold current is 20 mA.For a given inut current,  the output power measured is less compared to the value given in the data sheet. This is beacuse of the insertion loss of the isolator in the path.
  • The slope efficieny is calculted to be 0.018 mW/mA. The value given in the data sheet is 0.3 mW/mA. The reduction is because of the insertion loss of the isolator
  • The output of isolator is then connected to the AOM. Attachement #4 shows the schematic of the setup. The zeroth order port of the AOM is connected to the power meter and the output power from the first order port is blocked.
  • We then applied a dc voltage to the modulation input port of the AOM driver. We expect a decrease in the power from the zeroth order port with increase in the input voltage to the  AOM driver as the light is getting deflcted and power is geting coupled to the first order port.
  • Attachmet #5 shows the variation in power from zeroth order port with input voltage to the AOM driver.As expected, the power levels are decreasing with increase in input voltage
  • It is also observed that the output power levels are fluctuating. This could be because of temperature fluctuation and the thermal power meter is sensitive to that. 
Attachment 1: 2_micron_setup.jpg
2_micron_setup.jpg
Attachment 2: Laser_chara_setup.png
Laser_chara_setup.png
Attachment 3: Laser_characteristics.pdf
Laser_characteristics.pdf
Attachment 4: AOM_chara_setup.png
AOM_chara_setup.png
Attachment 5: AOM_characteristics.pdf
AOM_characteristics.pdf
  2322   Mon Apr 8 14:02:06 2019 AidanLab InfrastructureLab MonitoringHOBO MX100 temperature sensors/data loggers added to the optical tables

Here is the last 6 days.

Quote:

Here is the first set of data from the last 24 hours or so. The delta T  on the North table is around 5C versus 2C on the South table.

Quote:

I added a Bluetooth temperature sensor/data logger to each optical table. They're tucked INSIDE the NE corner of the enclosure structure - shielding them from the direct blast from the HVAC. They're recording the temperature once every 5 minutes. The came with double-sided tape to stick them into position.

The data can be accessed from your phone (I don't know yet if downloading the data clears the memory or not).

The idea is to gather data on the temperature of the two tables over the next few weeks.

Photos and instructions are attached.

 

 

Attachment 1: QIL_lab_temperature_6days.pdf
QIL_lab_temperature_6days.pdf
Attachment 2: QIL_South_2019-04-08_13_39_33_-0700.txt
Serial Number:	20572037 --

Date Time, GMT -0700	Temp, (*F)	Host Connect	EOF	
2019-04-02 14:05:45	78.25	 	 	
2019-04-02 14:10:45	76.47	 	 	
2019-04-02 14:15:45	71.84	 	 	
2019-04-02 14:20:45	70.68	 	 	
2019-04-02 14:25:45	70.37	 	 	
2019-04-02 14:30:45	70.30	 	 	
2019-04-02 14:35:45	70.37	 	 	
... 1720 more lines ...
Attachment 3: QIL_North_2019-04-08_13_39_16_-0700.txt
Serial Number:	20572038 --

Date Time, GMT -0700	Temp, (*F)	Host Connect	EOF	
2019-04-02 14:07:53	78.71	 	 	
2019-04-02 14:12:53	74.70	 	 	
2019-04-02 14:17:53	72.84	 	 	
2019-04-02 14:22:53	72.30	 	 	
2019-04-02 14:27:53	72.15	 	 	
2019-04-02 14:32:53	72.23	 	 	
2019-04-02 14:37:53	72.84	 	 	
... 1719 more lines ...
  2321   Wed Apr 3 12:51:55 2019 AidanLab InfrastructureLab MonitoringHOBO MX100 temperature sensors/data loggers added to the optical tables

Here is the first set of data from the last 24 hours or so. The delta T  on the North table is around 5C versus 2C on the South table.

Quote:

I added a Bluetooth temperature sensor/data logger to each optical table. They're tucked INSIDE the NE corner of the enclosure structure - shielding them from the direct blast from the HVAC. They're recording the temperature once every 5 minutes. The came with double-sided tape to stick them into position.

The data can be accessed from your phone (I don't know yet if downloading the data clears the memory or not).

The idea is to gather data on the temperature of the two tables over the next few weeks.

Photos and instructions are attached.

 

Attachment 1: QIL_North_2019-04-03_12_46_44_-0700.csv
Serial Number:20572038 --

"Date Time, GMT -0700","Temp, (*F)","Host Connect","EOF"
2019-04-02 14:07:53,78.71, , 
2019-04-02 14:12:53,74.70, , 
2019-04-02 14:17:53,72.84, , 
2019-04-02 14:22:53,72.30, , 
2019-04-02 14:27:53,72.15, , 
2019-04-02 14:32:53,72.23, , 
2019-04-02 14:37:53,72.84, , 
... 267 more lines ...
Attachment 2: QIL_South_2019-04-03_12_46_11_-0700.csv
Serial Number:20572037 --

"Date Time, GMT -0700","Temp, (*F)","Host Connect","EOF"
2019-04-02 14:05:45,78.25, , 
2019-04-02 14:10:45,76.47, , 
2019-04-02 14:15:45,71.84, , 
2019-04-02 14:20:45,70.68, , 
2019-04-02 14:25:45,70.37, , 
2019-04-02 14:30:45,70.30, , 
2019-04-02 14:35:45,70.37, , 
... 268 more lines ...
Attachment 3: QILtemperature.png
QILtemperature.png
  2320   Tue Apr 2 14:16:00 2019 AidanLab InfrastructureLab MonitoringHOBO MX100 temperature sensors/data loggers added to the optical tables

I added a Bluetooth temperature sensor/data logger to each optical table. They're tucked INSIDE the NE corner of the enclosure structure - shielding them from the direct blast from the HVAC. They're recording the temperature once every 5 minutes. The came with double-sided tape to stick them into position.

The data can be accessed from your phone (I don't know yet if downloading the data clears the memory or not).

The idea is to gather data on the temperature of the two tables over the next few weeks.

Photos and instructions are attached.

Attachment 1: IMG_9165.jpg
IMG_9165.jpg
Attachment 2: IMG_9158.jpg
IMG_9158.jpg
Attachment 3: IMG_9159.jpg
IMG_9159.jpg
Attachment 4: IMG_9160.jpg
IMG_9160.jpg
Attachment 5: IMG_9162.jpg
IMG_9162.jpg
Attachment 6: IMG_9163.jpg
IMG_9163.jpg
Attachment 7: IMG_9164.jpg
IMG_9164.jpg
  2319   Mon Apr 1 11:03:24 2019 JonUpdateComputingCymac assembly started

I know there is some CTN slow channel data on the disk. Is it at all possible to boot so that can be recovered?

Quote:

This afternoon Chris and I installed the ADC and DAC cards in fb4. We connected them to the timing card adapters (left external to the computer chassis for now).

We found fb4 to be running Debian 8 so first attempted to upgrade to 9, as that is the version supported by Jamie's cymac binaries. However, we encountered problems during the upgrade, apparently with gdm (the linux GUI). By switching to consol mode and killing gdm, were able to proceed to the point of updating all the packages. It completed successfully, but then the system failed to reboot, even in recovery mode. During boot, the advligo-rts kernel fails to start, and then boot hangs completely at the point the graphical interface is started.

We may want to start with a fresh install of Debian 9 and just reinstall the LIGO binaries.

 

  2318   Fri Mar 29 16:00:21 2019 awadeDailyProgressWOPORemeasuring TF Homodyne Photo Detector B after switching out foil cap

This set of measurments turned out to be bad. The OP27 (±15V supplies) with 50 Ω series at the output + 50 Ω load impedance seemed to be saturated.

The rest of this post is a stream of conciseness.

As noted in previous post (ATF:2316) the PD unit B in the homodyne had a slightly reduced HF response, possibly due to the use of a foil cap in the PD bias supply low pass.  I removed this and resmeasured.

Unit B TF: measured NF1611 DC voltage of  312 mV (@ 50Ω impedance), PD DC output voltage was - 566 mV (@ 50 Ω impedance).  Here the photodetector has transresistance is 6.8 kΩ. The NF1611 detectors have a DC path gain of 10 kΩ and an AC path gain of 700 Ω. From this the R_PC is xxx.  Here I ignore the overall phase (at ~1 MHz as the pi phase shift length is 150 m, so negligible).  This was a bad measurment, something was up with the circuit saturating way too low.  

Changed 100 Ω series on output to 50 Ω to make it more strait forward to understand how it works with 50 Ω load. 

Unit A TF: measured NF1611 DC voltage of  658 mV (@ 50Ω impedance), PD DC output voltage was 1.42 mV (@ 50 Ω impedance).  Here the photodetector has transresistance of 6.8 kΩ. The NF1611 detectors have a DC path gain of 10 kΩ and an AC path gain of 700 Ω. From this the R_PC is 0.3151.  Here I ignore the overall phase (at ~1 MHz as the pi phase shift length is 150 m, so negligible).  This measurment could have be saturating the current output of the OP27, likely that numbers here are not right.

After examining the output of detector B it became apparent that the output was saturated for powers of more than ~20 mA laser diode drive. 

Unit B TF: measured NF1611 DC voltage of  312 mV (@ 50Ω impedance), PD DC output voltage was -566 mV (@ 50 Ω impedance).  Here the photodetector has transresistance is 6.8 kΩ. The NF1611 detectors have a DC path gain of 10 kΩ and an AC path gain of 700 Ω. From this the R_PC is -0.3748.  Here I ignore the overall phase (at ~1 MHz as the pi phase shift length is 150 m, so negligible).   100 Ω  (50Ω + 50 Ω) loading is too low, OP27 open loop gain and voltage swing range slew starts to sag below 150Ω loading.  These results are bogus.

This time I measured with 50 Ω on the oscilloscope for apples-to-apples comparison with what the Aglient 4395A was seeing.  Turns out I think I was saturating the photodetector current wise in the first round of measurments. All of the above is probably not right and should be disregarded.  Also, later I switched to a AD829 as the op amp of choice so the above is not much use for characterizing the detector performance anymore.  The OP27 isn't really spec'ed to operate with loads this low and only has its full band with for output loads of >2 kΩ.  

  2317   Thu Mar 28 18:16:59 2019 JonUpdateComputingCymac assembly started

This afternoon Chris and I installed the ADC and DAC cards in fb4.

  • We connected them to the timing card adapters (left external to the computer chassis for now).
  • We found fb4 to be running Debian 8 so first attempted to upgrade to 9, as that is the version supported by Jamie's cymac binaries.
  • However, we encountered problems during the upgrade, apparently with gdm (the linux GUI). 
  • By switching to consol mode and killing gdm, were able to proceed to the point of updating all the packages.
  • It completed successfully, but then the system failed to reboot, even in recovery mode.
  • During boot, the advligo-rts kernel fails to start, and then boot hangs completely at the point the graphical interface is started.

We may want to start with a fresh install of Debian 9 and just reinstall the LIGO binaries.

  2316   Wed Mar 27 20:37:38 2019 awadeDailyProgressWOPOMeasuring TF Homodyne Photo Detector for WOPO experiment

Initial measurment of PD TF

I realize I never really measured the signal transfer function for each of these Photo Detectors (PD).  This post summarizes measurements of the optical to electrical signal output signal transfer function. Here I used the Jenne rig at the 40m and took a transfer function using an Aglent 4395A from 30 kHz to 5 MHz.  

I've labeled the two PD in the homodyne as unit A and unit B.  Below (first attached)  is the raw TF as measured by the Aglient taking the ratio of the detector to the reference NF1611 detector:

Fixing differences between output impedance unit A and B

After looking a bit at the circuit I realized that I had added 100 Ω in series in detector B (to limit current draw when driving 50 Ω loads).  This had not been added to the detector A, which means that with more power than 300 µW on this detector the op amp would be drawing more than 30 mA when driving a 50 Ω.  This shouldn't have been an issue when driving the summing circuit, but is good practice.  

I added 100 Ω in series with op27 output (thin film 1206 size) in detector A to match the output to that of detector B.  I remeasured the TF and get a much better match between the paths.

 

Calibrating TIA Measurment

The following is used to calibrate the  transfer function (See PSL:2247):

Z_\textrm{AC,PD} = Z_\textrm{AC,Ref} R_\textrm{PC}e^{i\delta\phi} T_\textrm{meas}

where ZAC,PD is Calculated RF Transimpedance, Z AC,Ref is the known RF transimpedance of the reference photodiode, 𝛿ϕ is arbitrary phase delay due to light and cable length and RPC  is the photocurrent ratio at DC of reference PD to RFPD under test given by

R_\textrm{PC} = \frac{V_\textrm{DC,Ref}/Z_\textrm{DC,Ref}}{V_\textrm{DC,PD}/Z_\textrm{DC,PD}}

For these measurements the calibration factors were as follows:

Unit A TF (after adding 100 Ω series​ output): measured NF1611 DC voltage of 780 mV (@ 1 MΩ impedance), PD DC output voltage was -1.20 V (@ 1 MΩ impedance).  Here the photodetector has transresistance is 6.8 kΩ. The NF1611 detectors have a DC path gain of 10 kΩ and an AC path gain of 700 Ω. From this the R_PC is 0.4496.  Here I ignore the overall phase (at ~1 MHz as the pi phase shift length is 150 m, so negligible).  

Unit B TF: measured NF1611 DC voltage of 791 mV (@ 1 MΩ impedance), PD DC output voltage was -1.18 V (@ 1 MΩ impedance).  Here the photodetector has transresistance is 6.8 kΩ. The NF1611 detectors have a DC path gain of 10 kΩ and an AC path gain of 700 Ω. From this the R_PC is 0.4558.  Here I ignore the overall phase (at ~1 MHz as the pi phase shift length is 150 m, so negligible).  

 

Results of calibrated transimpedance measurements are plotted below*. I also added the LISO estimated noise curve.  The 3 dB of unit A and B were 1.93 MHz and 1.72 MHz respectively.  There is some discrepancy with the LISO model here, some of this might be artifacts in the reference detector and some is likely dirt effects in the proto board circuit that I'm not going to debug for now.  The biggest discrepancy between the two detectors starts around 900 kHz.  When I looked again at the detectors I realized I had used a Wima foil cap in detector B for the bias LPF. It should be fine for these frequencies, but I switched the foil cap out and replaced it with a 1 µF + 100 nF ceramic cap to match the unit A.  I haven't gotten a chance to remeasure this since as Gautam is still using the Jenne rig to do some stuff. I'll remeasure later, we should then find the two responses match.

 

*I didn't do the many IRIS measurement​ uncertainty analysis, its not necessary at this stage. 

 

Attachment 1: TFUnitAandB_BeforeMatchingOutImp.pdf
TFUnitAandB_BeforeMatchingOutImp.pdf
Attachment 2: TFUnitAandB_FixingDetectorAImp.pdf
TFUnitAandB_FixingDetectorAImp.pdf
Attachment 3: WOPO_HD_TranImpGain_UnitAandBandLISO.pdf
WOPO_HD_TranImpGain_UnitAandBandLISO.pdf
  2314   Wed Mar 27 11:27:00 2019 AidanLab InfrastructureHVACHVAC temperature gradients in the QIL

Someone from Facilities named Modesto came out to the QIL to see the problem with the HVAC. He measured 80F coming out of the hot vent (when its thermostat was set to 68F) and 65F coming out of the cold vent. He was able to drop the hot vent to about 62F by turning the set point down to 55F (so the controller is at least working).

I asked that he set it up that both units are controlled by one temperature sensor. So he’s going to disconnect the hot vent controller from it’s thermostat and tie it into the thermostat for the cold side. This seems reasonable but could be potentially problematic if the HVAC units have significantly different responses.

Anyway, the current plan is for him to come in on Friday to do this work

  2313   Tue Mar 26 17:13:53 2019 AnjaliUpdate2micronLasersFrequency stabilization of 2 micron source

Experimental plan 

  • The aim is to repeat the homodyne measurement setup with  balanced detection scheme for the frequency stabilisation of 2 micron laser source.
  • Attachment # 1 shows the schematic of the experimental setup.
  • All the components specified in the schematic are available.
  • For the balanced detection, we can initialy use SR560, but we need to design a new circuitry for the same.
Attachment 1: Homodyne_setup_for_2_micron.pdf
Homodyne_setup_for_2_micron.pdf
  2312   Mon Mar 25 09:40:16 2019 AnjaliNoise Budget2micronLasersNoise budget for the frequency discriminator
  • Different fiber parameters are very important in determining the thermal noise of the fiber. I was considering SM1950 fiber for the calculations, but all the characteristic parameters for this fiber are not known.
  • Glenn ( IEEE Journal of Quantum Electronics25(6), 1218-1224 ) and Jing Dong (Applied Physics Letters108(2), 021108) had given the set of fiber parameters. Attachment #1 shows the comparison of thermoconductive noise of the fiber with fiber parameters taken from Glenn and Jing Dong. The parameters from Jing Dong are for Corning standard SMF 28 fibers and they are the recent resuts compared to Glenn. Hence I think it would be good to follow the parameters given in Jing Dong's paper.
  • It is also found that the thermal noise of fiber has contributions from the thermoconductive noise (https://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=255919) and thermomechanical noise (https://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=5614023) . Thermoconductive noise  (in rad/rtHz) predicts a frequency independent characteristics at lower frequencies where as thermomechanical noise (in rad/rtHz) has a 1/f characteristics over a certain frequency range. The thermomechanical noise is of concern only over a certain frequency range it is given by the ratio of velocity of sound with the total length of the fiber. This indicates that, in the case of thermomechanical noise, the fluctuation of concern is much slower than the time it takes for the sound wave to travel across the length of the fiber. 
  • Attachment #2 shows the noise budget including the thermomechanical noise of the fiber. With 10 m length of the fiber, the thermoconductive noise is of concern for frequencies lower than 364 Hz . Thus a low pass transfer function is assumed with corner frquency of 364 Hz in the calculation of thermomechanical noise. 
Attachment 1: Comparison_of_thermoconductive_noise_in_Hz_per_rtHz.pdf
Comparison_of_thermoconductive_noise_in_Hz_per_rtHz.pdf
Attachment 2: Noise_budget_with_thermomechanical_noise.pdf
Noise_budget_with_thermomechanical_noise.pdf
  2311   Fri Mar 22 10:00:30 2019 AnjaliNoise Budget2micronLasersNoise budget for the frequency discriminator

Attachement #1 shows the modified noise budget with modification on dark current noise.

I still have to find out the PSD for room temperature to plot the room thermal noise

 

Attachment 1: Noise_budget_modified.pdf
Noise_budget_modified.pdf
  2310   Thu Mar 21 09:43:19 2019 AnjaliNoise Budget2micronLasersNoise budget for the frequency discriminator

Also, I have a doubt that whether the transfer function should be of low pass in nature. So, if the laser phase is fluctuating faster than the time it takes to propagate throught the delay fiber, we will not be able to discriminate .

If that is the case, the noise budget looks like as shown in attachment #1

Attachment 1: Noise_budget_with_low_pass_transfer_function.pdf
Noise_budget_with_low_pass_transfer_function.pdf
  2309   Thu Mar 21 00:49:35 2019 AnjaliNoise Budget2micronLasersNoise budget for the frequency discriminator

I didn't use the transfer funtion in the calculation of dark current noise thinking dark current noise is a characteristic of the detector alone

Regarding room thermal fluctuation, I was plotting the thermal noise of the detector which is given in A/rtHz as 

=\sqrt{\frac{4k_bT}{R_f}}

  2308   Wed Mar 20 19:17:50 2019 ranaNoise Budget2micronLasersNoise budget for the frequency discriminator

I think the dark noise should be very close to the shot noise and also have the same transfer function.

Also, the room temperature fluctuations are probably large at low frequencies. What PSD for room temperature did you use?

 

  2307   Wed Mar 20 15:57:55 2019 AnjaliNoise Budget2micronLasersNoise budget for the frequency discriminator

I was trying to prepare the noise budget for the frequency stabilisation setup for 2 micron laser. In this test, we use a fiber based Mach-Zehnder interferometer as a frequency discriminator to convert frequency noise to amplitude(voltage) fluctuations. The different noise sources to be considered in this analysis are the following

  1. Shot noise
  2. Dark current noise
  3. Room thermal noise
  4. Fiber thermal noise
  5. Fiber acoustic noise
  6. Laser intensity noise
  7. Photothermal noise
  • Attachment #1 shows the magnitude spectrum of the transfer fucntion for the frequency discriminator. I assumed the delay fiber length in one of the arms of the Mach-Zehnder interferometer is 10 m. This corresponds to a time delay of 50 ns. The corner frequency is inversely proportional to the time delay.
  • The transfer fucntion can be interpreted as, if the frequency variation is very slow- it can't be distiguished by the discriminator and we dont get any signal.
  • Attachment # 2 and # 3 shows the shot noise and fiber thermal noise respectively in rad/rtHz
  • The shot noise is calculated  considering thorlabs DET10D detector with a responsivity of 1.2 A/W at 2 micron. The laser output power from Eblana laser is assumed to be 2 mW and the power reaching the detector is estimated to be 1.23 mW.
  • The fiber thermal noise is calculated based on Wanser, K. H. (1992). Fundamental phase noise limit in optical fibres due to temperature fluctuations. Electronics letters28(1), 53-54.
  • The dark current noise and room thermal noise are calculted to be a constant value (3.97x10-12 rad/rtHz and 1.20x10-12 rad/rtHz respectively)- but this is not correct
  • The noise in rad/rtHz is converted to Hz/rtHz by mutiplying with the frequency (f) and it is shown in attachment #4
Attachment 1: High_pass_transfer_function.pdf
High_pass_transfer_function.pdf
Attachment 2: Shot_noise.pdf
Shot_noise.pdf
Attachment 3: Thermoconductive_noise_of_fiber_Glenn_parameter.pdf
Thermoconductive_noise_of_fiber_Glenn_parameter.pdf
Attachment 4: noise_budget_combined_log_log.png
noise_budget_combined_log_log.png
  2306   Wed Mar 13 17:49:14 2019 awadeSummaryWOPODamage threshold estimates for optical fibers

This is just a note about damage tolerances for fibers so we have a reference of the amount of power that can be used.  

Thorlabs gives the maximum theoretical CW power as 1 MW/cm² and a 'practical' safe power level of 250 kW/cm² (or a quarter of the max).  They don't seem to provide information about wavelength dependence and assume its the same for all wavelengths.  We don't expect to be affected by any of the exotic ultra high power effects like bend loss induced damage and photodarkening.  The air/glass interface is where the damage will occur.  This can either be because of heating of the ferrule/connector (causing epoxies etc to break down and damage the interface by depositing on the optical surface) or regular mechanism that are the same as bulk optics (dielectric break down and thermal effects).  

The intensity profile of light confined in the fiber is defined by the Mode Field Diameter (MFD) -- the cross-sectional diameter of the light that includes the core of the fiber and a region just beyond the cladding the mode occupies. MFD of 1064 PM fiber (PM980-XP) is 6.6 ± 0.5 μm @ 980 nm and for the 532 nm PM fiber (PM460-HP) is 3.3 ± 0.5 µm @ 515 nm.  

Fiber effective area is 

A = \pi (\textrm{MDF}/2)^2

which is 8.5x10^-8 cm^2 for PM460-HP and 3.4x10^-7 cm^2 for PM980-XP.  Taking the conservative 'practical' damage threshold this indicates a maximum power of 21.4 mW into the 532 nm fiber and 85.5 mW into the 1064 nm PM fiber. The absolute maximum is just a factor of four more than this: 85.6 mW into 532 nm PM fiber and 342 mW into 1064 nm fiber. If the fiber ends are kept clean then we should be fine if the power level is kept below 85 mW.

 

Fiber beam splitter

Of particular concern is the power handling capability of the fiber beam splitter (PN1064R5A2).  There will be some waste 532 nm light coming out of the WOPO and I don't want these potentially multi-mode components to exit the cladding at the point of the coupler and damage the surrounding material. The 1064 nm maximum power rating of this 50:50 PM beam splitter is listed as 1 W (for the connectorized fiber), so we should be well clear of that threshold for the LO light.  For 532 nm its less clear. The equivalent 532 nm PM 50:50 beam splitter (PN530R5A2) has a rated power of 100 mW @ 530 nm for bare or connectorized fibers. As the MFD of the 1064 nm version of this PM beam splitter has a much larger MFD and the exiting 532 nm light will already be expanded in the cable patching the WOPO to the BS, we should be well clear of this damage threshold point.  

So bottom line is that we need to keep power below 85 mW going into the WOPO device and keep all the end connectors super clean and it will be fine.

 

 

  2305   Wed Mar 13 12:44:41 2019 awade, anchalDailyProgressWOPOPol launch into PM fiber 1064 nm

[awade, anchal ]

After a bit of reading I've realized that the standard use of these PM fibers is to launch along the slow axis (see for example Thorlabs and OzOptics info on fiber beam splitters).  It should be much of the sameness for patch cables, but polarization sensitive elements like beam splitters are mostly tested and specified for slow axis launch unless they are custom made to order. 

We are switching the polarization alignment to slow axis in the 1064 nm and 532 nm fiber coupling.  Anchal is re-optimizing​ the 1064 nm launch to get the PM fiber extinction ratio back to a good place.  We've also changed input launch to use a laser line PBS mounted in a rotation mount for clean linear polarization.  With the optimized setup the for the 1064 nm fiber path the output polarization signal goes from 3700 mV to 39.3 mV which is an extinction​ ratio of -19.7 dB.

Here the max theoretical extinction​ ratio is 

ER = -10 \log_{10}[\tan^2(\theta)]

which would place our goodness of alignment to with 0.61 deg.

Updated 1064 nm launch. Uses rotation mounted PBS for guaranteed linear
polarization, half wave plate is to maximize power.

 

 

Quote:

I've replaced the SM fiber in the 1064 nm launch with a PM fiber (P3-1064PM-FC-5). I also moved the fiber collimator (F240APC-1064) back 2.54 cm back to give more space for a PBS cube (to check linearly of the light).  

For the 1064 nm launch it seemed to be a lot harder to find the initial alignment of the collimator using the alignment of the back propagated 650 nm fiber laser source. Here I aligned a pair of irises in the forward propagating direction and then back propagated through the PM fiber using 650 nm to get the initial​ pointing of collimator. I don't know why this is so much harder than the 532 nm case.  I suspect one of the steering mirrors is not really reflecting off the front dielectric surface.  In the end I did a bunch of systematic walking of the fiber launch mount and eventually fount the alignment.  

From 4.44 mW of input light I get 2.74 mW of light out the other end of the fiber.  This is an efficiency of 62 % which is more than enough for my needs.  I expect the HD will only need 1 mW (2 mW max), so this is fine. Getting this in coupling higher will require a bit of lens walking, not really worth it at this stage.

I had already carefully aligned the collimator orientation to put the fast axis on aligned to p-pol (wrt the table), by eye.  It seems like the launch pretty much hit the correct launch polarization on the first go.  I see little variation in the polarization when I pulse the heat on the fiber.  This is now good to go for optimizing the homodyne visible and polarization overlap output from the SQZ.

 

  2304   Tue Mar 12 16:58:50 2019 awadeMiscWOPOTodo list W11

These are things to get do this week on WOPO experiment:

☑️ Reinstall 50:50 fiber splitter into homodyne setup and go back to fiber launch of both ends of the HD directly onto photodetectors (rather than free space)

☐ Check visibility of HD by launching 1064 nm into both arms of HD using splitter and extra length on one arm, ramp laser frequency to get fringes and look at Vpp on each diode seperatly

☑️ Optimize subtraction of HD for max Common Mode Rejection (CMR) of LO amplitude noise.  Inject 3.21 kHz line into laser BNC port and minimize this peak on the subtracted output of the HD.

☐ Check 1064 nm -> 532 nm conversion in WOPO device to establish polarization basis for correct pol alignment into fiber (change 532 nm launch polarization if necessary), should be along the fast axis but for some reason this isn't in the datasheet explictly

☐ Double check how much power change there is from 2 pi modulation depth from PZT mounted 1064 nm mirror.  We don't want to over actuate on this element as it slightly misaignes into the fiber launch and causes some change in power, this is ok over a small range as the CMR of the homodyne will reject this.  We just want to be sure that this isn't the dominant effect that we are seeing at the output once we thing we should be seeing SQZ.

 

☑️Tick mark when done.

  2303   Tue Mar 12 16:35:43 2019 awadeDailyProgressWOPOPol launch into PM fiber 1064 nm

I've replaced the SM fiber in the 1064 nm launch with a PM fiber (P3-1064PM-FC-5). I also moved the fiber collimator (F240APC-1064) back 2.54 cm back to give more space for a PBS cube (to check linearly of the light).  

For the 1064 nm launch it seemed to be a lot harder to find the initial alignment of the collimator using the alignment of the back propagated 650 nm fiber laser source. Here I aligned a pair of irises in the forward propagating direction and then back propagated through the PM fiber using 650 nm to get the initial​ pointing of collimator. I don't know why this is so much harder than the 532 nm case.  I suspect one of the steering mirrors is not really reflecting off the front dielectric surface.  In the end I did a bunch of systematic walking of the fiber launch mount and eventually fount the alignment.  

From 4.44 mW of input light I get 2.74 mW of light out the other end of the fiber.  This is an efficiency of 62 % which is more than enough for my needs.  I expect the HD will only need 1 mW (2 mW max), so this is fine. Getting this in coupling higher will require a bit of lens walking, not really worth it at this stage.

I had already carefully aligned the collimator orientation to put the fast axis on aligned to p-pol (wrt the table), by eye.  It seems like the launch pretty much hit the correct launch polarization on the first go.  I see little variation in the polarization when I pulse the heat on the fiber.  This is now good to go for optimizing the homodyne visible and polarization overlap output from the SQZ.

  2302   Mon Mar 11 08:49:11 2019 Aidan, ChubComputingCymacsBuilding FB4 Cymacs

We started migrating equipment for the FB4 Cymacs to the QIL on Friday. See attached list and images.

Attachment 1: IMG_8851.jpg
IMG_8851.jpg
Attachment 2: IMG_8852.jpg
IMG_8852.jpg
  2301   Fri Mar 8 17:05:52 2019 AnjaliUpdate Frequency stabilization of 2 micron source

 

The schematic of the homodyne configuration is shown below.

Following is the component list

Item Quantity Availability Part number Remarks
Laser diode 1 Yes EP 2004-0-DM-B06-FA  
Isolator 1 Yes    
Coupler(50/50) 5 Yes TW2000R5A2A Fiber type : SM2000
Delay fiber   Yes SM2000 Need to check the exact length of the fiber
Photodetector 4 Yes DET10D  
SR560 3 Yes    

 

Attachment 1: Homodyne_setup_2micron.png
Homodyne_setup_2micron.png
  2300   Thu Feb 28 16:00:59 2019 awadeDailyProgressWOPOPol launch into PM fiber

I've set up a rotating PBS and half-wave plate to provide polarization adjustment into the 532 nm fiber without misalignment the spatial alignment.  Here I've used a PRM1 rotation mount with a SM1PM10 lens tube mount for beam cube prisms.  The lens tube mount is supposed to be for pre-mounted cubes but I've inserted some shims to hold it in place and it seems to work well like that.  It means I can get a nice clean linear polarization at all rotations.

After spatially aligning the input beam I stepped the rotation of the PBS (and accordingly the L/2 wave plate) and pulsed the temperature of the fiber using a heat gun.  After some walking I found that for the current fiber rotation (0 deg) the linear polarization was aligned with the fiber axis at 88 deg PBS rotation (here 0 deg PBS rotation is aligned for p-pol transmission, well almost). I made some adjustments to the alignment of the fiber collimator in the fiber launch, I aligned the slow axis key with the vertical so that the fast axis of the fiber is p-pol.

Keying on PM fibers

As a side note the keying of PM fiber patches is typically with the slow axis aligned with the key notch. The WOPO's PM fibers are keyed so that the alignment key is along the slow axis of the fiber (i.e. aligned with the stress rods). Figure below illustrates the configuration.

Replacing the 532 nm patch with fresh PM fiber

I was getting a large jitter in the power levels as measured at the output of the old SM and PM fibers (on the order of 10%).  These power fluctuations were not present on the input side.  I thought this was an alignment jitter or a polarization effect.  However, I was unable to minimize it by improving the input polarization at the launch.  When I tapped various mounts there didn't seem to be a corresponding correlation with output power jitter of the fiber.  When I checked the end of the PM fiber (P3-1064PM-FC-2), I saw that there was damage about the core (see pictured below).  It seems like maybe I had some kind of etalon effect from this burn mark and the launch.  After replacing the 532 nm PM fiber with a fresh one that arrived last week the power is much more stable and I was able to easily​ find the pol alignment going in. 

Damage to PM fiber end.  No amount of aggressive cleaning will remove the mark in the middle.
The fiber will need to be cut and a new connector spliced on.

 

Next job is to replace SM fiber for the 1064 nm delivery with PM fiber so there is a well defined polarization for launching into the homodyne detector.

 

 

Quote:

Alignment of the pumping 532 nm polarization into the WOPO is important to getting the correct phase matching condition.  For the periodically polled Lithium Niobate (LN) waveguide the phase matching is type-0: and pumping and fundamental wavelengths are in the same polarization.  The AdvR non-linear device is coupled with polarization maintaining fibers (Panda style), which are keyed at their FC/APC ends.  This means that with the correct launch polarization we should be correctly aligned with the proper crystal axis for degenerate down conversion (at the right chip temperature). 

Replacing Broadband PBS

Till now I was using non-pol maintaining patches to coupling into the WOPO fiber ends.  This should have been ok, but it is hard to figure out exactly which polarization is optimal so I switched to a pol-maintaining patch because it can be aligned separately and then the keyed connectors give you automatic alignment.  I had some issues trying to find the optimal polarization going into the fiber and I've now traced this back to the polarizing beam cubes.  I've been using Thorlabs PBS101 which is a 10x10x10 mm^3 beam cube that is supposed to be broad-band (420-680 nm).  When I checked the extinction ratio I saw Pmax=150 mW, Pmin=0.413 mW on transmission between extremes.  This is an extinction ratio of Tp:Ts = 393:1 which is much less than the spec of >1000:1.  Not sure what's going on here, the light going into the BS is coming directly from a Faraday isolator and a half-wave plate.  With some adjustment to the angle of the wave plate I can do a little better but it should be nicely linearly polarized to start with.

I've switched out the PSB101 for the laser line PBS12-1064 I remeasured extinction ratio (Pmax=150 mW, Pmin=27.6 µW) Tp:Ts = 5471:1 (better than the quoted 3000:1 spec).  This is good, at least now I know what is going on. I am also putting in an order for a 532 nm zero order quarter-wave plate, so that we can be absolutely sure we are launching in linear light always.  

Aligning light into pol-maintaining fiber

I previously thought I might be able to use the frequency modulation technique to align the light through the polarization maintaining fiber.  There is a birefringence in PM460-HP fiber of  3.5 x 10-4.  The phase between ordinary and extraordinary axes over the whole fiber length is

\Delta\phi = \frac{2 \pi \Delta n L}{c}f

Where L is fiber length, \Delta n is the birefringence and f is the laser frequency.  The idea is to launch linearly polarized light into the fiber and then at the readout place a polarizer rotated to be 90°: ramping frequency will produce an amplitude modulation on the dark fringe.  However, even with 1 GHz of frequency ramp this is only a 15 mrad effect for a 2 m fiber, its likely to be too small to see over other effects.  This is not enough to be able to fine align polarization.  

Instead I'll use the heat gun method.  I'll fire linearly polarized light into the fiber and measure the output with a crossed polarizer.  If the input polarization is correct there should be no power changes on the output as the fiber is thermally cycled. Its only two meters long so hopefully this effect is easy to see.

 

Attachment 2: PMFiber.pdf
PMFiber.pdf
  2299   Mon Feb 25 20:42:32 2019 KojiSummaryScatterometerSi Block stored in Cryo Lab

On the Friday cleaning, we vacated the east optical table. The Si scatterometer was disassembled and the Si block was moved and stored to the cryo lab.

Attachment 1: P_20190221_115501.jpg
P_20190221_115501.jpg
  2298   Fri Feb 22 14:05:09 2019 awadeDailyProgressWOPOMixing down homodyne: 5th order elliptic filter (100 kHz)

Making 5th order Elliptical Filter

I couldn't find any filters that would cut off above 100 kHz so I made my own using a Thorlabs EEAPCB1 generic filter PCB in a Thorlabs EEA14 enclosure. I used a 5th order elliptic design with a pass band up to 100 kHz and a stop band of 40 dB from 150 kHz.  To speed things up I used the Coilcraft Filter Designer software  (4.0.1, Windows) and chose closest standard values from parts we had in EE workshop kits. The Coilcraft designer is nice because it has the full physical model of the inductors built in.

The schematic is illustrated below:

5th order Elliptic Filter For use in homodyne readout electronics.  These are ideal values, actual values closest values were chosen from capacitor and inductor kits (listed in main text)

Actual values selected were closest available and I didn't try to do any mixing or matching to get fine tuned correct values. Values are as follows:

  • L2 and L4 27 µH => 1812CS-273XJL (27µH @ 2.5 MHz, tol 5%, Q_min 15 @ 2.5 MHz, SRF_min 11 MHz, DCR 11.04 Ω)
  • C1 20.47 nF => 12065C223KAT2A (0.022 µF, 50 V, 10%)
  • C2 14.38 nF => 12065153KAT2A (0.015 µF, 50 V, 10%)
  • C3 47.42 nF => 12061C473KAT2A (0.047 µF, 50 V, 10%)
  • C4 4.792 nF => 12065C472KAT2A (4700 pF, 50 V, 10%)
  • C5 27.3 nF => 12065C273KAT2A (0.027µF, 50 V, 10%)

The built filter is shown below.

The transfer function was taken from 10 Hz up to 5 MHz (IF BW of 10 Hz).  The Coilcraft Filter Designer software seems to export all the filter scattering parameters except S12 and S21, not the best. Instead I used LISO to model the filter's predicted response and this is plotted alongside measured TF.  Here I have scaled the filter model's response by 2 to match the impedance condition of the TF measurement.  The coilcraft 0805LS-273X_E 27 µH inductors series resistance was modeled as 11 Ω (as per spec sheet) and values of capacitors were those used in the actual circuit.  In the original ideal 5th order elliptical circuit there is a double dip above the corner frequency.  The series resistance of the non-ideal inductors dampens these.  I don't really want to spend much more time on mixing and matching capacitor values.  It looks like for now that the pass band ripple is acceptable and the attenuation is >40 dB at 1 MHz and 2 MHz where we are trying to block harmonic signals. I'll leave optimization of this part for now and write this off as done.

Edit Mon Feb 25 19:22:14 2019 (awade): Fixed the phase pane of the bode plot, had accidentally used magnitude and wrong units. Also fixed some spelling.

Attachment 3: Elliptical5thOrder_MeasuredVsModel.pdf
Elliptical5thOrder_MeasuredVsModel.pdf
Attachment 4: 20190222_5thOrderEllipticFilter100kHz.zip
  2297   Thu Feb 21 17:02:24 2019 awadeDailyProgressWOPOMixing down homodyne output from 1 MHz

I want to get signal from about 1 MHz down to around DC from my subtracted  homodyne photodetectors. I'm planning to do something like this:

Level 7 mixer with low pass (1.9 MHz) filter + 50 Ω terminator on output. LO will come from a
RF function generator and R port will come directly from photodetector subtraction.

This mixer has a 4.78 dB conversion loss and should do the job.  Only issue is that to operate the mixing down the low pass filter needs to be lowered from 1.9 MHz down to something pretty low to ensure the harmonic (2 MHz term is removed).  These are 4th order filters, we'd probably want the cut off to be an order of ten below the mixing frequency... 100 kHz.  I don't see this in the minicircuits catalog don't know how doable that is to make one.  I'll have a look at what the 40 m has.

Attachment 1: IMG_5106.JPG
IMG_5106.JPG
  2296   Wed Feb 20 10:38:46 2019 awadeDailyProgressWOPOFixing 532 nm polarization linearity issues

Alignment of the pumping 532 nm polarization into the WOPO is important to getting the correct phase matching condition.  For the periodically polled Lithium Niobate (LN) waveguide the phase matching is type-0: and pumping and fundamental wavelengths are in the same polarization.  The AdvR non-linear device is coupled with polarization maintaining fibers (Panda style), which are keyed at their FC/APC ends.  This means that with the correct launch polarization we should be correctly aligned with the proper crystal axis for degenerate down conversion (at the right chip temperature). 

Replacing Broadband PBS

Till now I was using non-pol maintaining patches to coupling into the WOPO fiber ends.  This should have been ok, but it is hard to figure out exactly which polarization is optimal so I switched to a pol-maintaining patch because it can be aligned separately and then the keyed connectors give you automatic alignment.  I had some issues trying to find the optimal polarization going into the fiber and I've now traced this back to the polarizing beam cubes.  I've been using Thorlabs PBS101 which is a 10x10x10 mm^3 beam cube that is supposed to be broad-band (420-680 nm).  When I checked the extinction ratio I saw Pmax=150 mW, Pmin=0.413 mW on transmission between extremes.  This is an extinction ratio of Tp:Ts = 393:1 which is much less than the spec of >1000:1.  Not sure what's going on here, the light going into the BS is coming directly from a Faraday isolator and a half-wave plate.  With some adjustment to the angle of the wave plate I can do a little better but it should be nicely linearly polarized to start with.

I've switched out the PSB101 for the laser line PBS12-1064 I remeasured extinction ratio (Pmax=150 mW, Pmin=27.6 µW) Tp:Ts = 5471:1 (better than the quoted 3000:1 spec).  This is good, at least now I know what is going on. I am also putting in an order for a 532 nm zero order quarter-wave plate, so that we can be absolutely sure we are launching in linear light always.  

Aligning light into pol-maintaining fiber

I previously thought I might be able to use the frequency modulation technique to align the light through the polarization maintaining fiber.  There is a birefringence in PM460-HP fiber of  3.5 x 10-4.  The phase between ordinary and extraordinary axes over the whole fiber length is

\Delta\phi = \frac{2 \pi \Delta n L}{c}f

Where L is fiber length, \Delta n is the birefringence and f is the laser frequency.  The idea is to launch linearly polarized light into the fiber and then at the readout place a polarizer rotated to be 90°: ramping frequency will produce an amplitude modulation on the dark fringe.  However, even with 1 GHz of frequency ramp this is only a 15 mrad effect for a 2 m fiber, its likely to be too small to see over other effects.  This is not enough to be able to fine align polarization.  

Instead I'll use the heat gun method.  I'll fire linearly polarized light into the fiber and measure the output with a crossed polarizer.  If the input polarization is correct there should be no power changes on the output as the fiber is thermally cycled. Its only two meters long so hopefully this effect is easy to see.

  2295   Tue Jan 8 11:38:49 2019 awadeLab InfrastructureGeneralRoughing pump leaking oil on bake rig

The roughing pump attached to the bake rig in the QIL (room B265B) is leaking oil.  It seems to be coming out of the box that should be filtering the exhaust (pictured below).

Its been clean up but please don't use bake rig for now and take care when entering lab.

Oil seems to be coming out of the exhast filter

I'll ask Chub to have a look.

Edit Tue Jan 8 19:58:03 2019 (awade): Chub has now fixed this. He cleaned the filter and installed a breather tube with a cloth wire-tied to the end to prevent futher spillage

Attachment 2: IMG_4862.JPG
IMG_4862.JPG
  2294   Mon Jan 7 19:27:10 2019 awadeDailyProgressHomodyneFixing homodyne fringe visibility and estimating losses

Fringe visibility of the homodyne was indeed not great.

It turns out that the combination of poor spatial overlap, polarization overlap and mode matching I couldn't see any fringes formed between the two arms of of the homodyne (when they were excited from the same laser source). 

Fixing fringe visibility

To improve spacial and polarization overlap between the two fiberized inputs I need to measure the fringe visibility, i.e. peak to peak fringe contrast. The inputs of the homodyne were connected a 50/50 fiber splitter, an extra 1 m of patch cable was used in the signal arm as follows:

Frequency modulated 1064 nm laser light is launched into both arms of homodyne
detector.  Here the signal arm patch cable is 1 m longer than the other arm giving the 
single PD FSR of 300 MHz.  The peak to peak fringes give the visablity.

With the extra length in the signal arm we have a Mach–Zehnder with an FSR  of 300 MHz.  For the Diabolo the NPRO PZT has a quoted response of 1.95 MHz/V, so 77 V is enough to see a full peak-to-peak of the fringes.

After some initial alignment of the two input beams by eye with a slow scan of the laser frequency, I was able ot see some fringes on the combined beams.  Looking at the signal on a single detector is was possible to walk the alignment between the two paths until the Vmax = 4.14 V, Vmin = 1.6 V.  Here fringe visibility is:

\textrm{Fringe Vis.} = \frac{V_\textrm{max}-V_\textrm{min}}{V_\textrm{max}+V_\textrm{min}}

With initial positioning of the fiber collimator launches I got 44 % visibility.  The quantum efficiency of the overlap scales as the square of the visibility (\eta_\textrm{h}=V^2).  Thus based on the initial placement of the fiber launches the maximum inferred quantum efficiency would be 0.19, not great.  

On closer inspection of the interference pattern, it was apparent that the interference was forming a bull's eye pattern.  The propagation​ distance from the signal arm was 20 mm shorter than that of the LO arm.  I moved the fiber launch back to match the arm lengths and realigned the beams.  Now with 125 mm distance for both arms to the beam splitter I was able to optimize to Vmax = 4.9 V and Vmin = 0.320 V, giving a visibility of 88%. That means a minimum quantum efficiency due to overlap of 0.77.  This could be an underestimate of the visibility and efficiency if the polarizations were not optimal.  More likely the power ballance between the arms would have biased the measurement.  I didn't note down the power measured from each of the launches, from memory it was to within 5%.

Another thing to note is that the patch cables used in the launch are slightly different.  Even though the collimators are both F240APC-1064 the slightly different MFD of the two fibers means that the launched beam will have a slightly different waist (position​ and radius).  Maybe mode matching could be improved with the purchase of a matching patch for the signal path to the LO path.

Total estimates of losses (Worst case)

Total estimated losses are max 0.2 dB/cm within the waveguide (total max .91), 0.5 at the fiber WOPO butt coupled interface, 0.7 dB (typical/max?) insertion loss at fiber to fiber interface x1 (~ 0.15), 0.23 due to mode overlap loss at HD, 0.15 at the photodiode themselves.  This would total 0.25 quantum efficiency​.  From this the lower bound of max squeezing would be

 SQZ_\textrm{Max} = 10 \log[1-\eta_\textrm{total}]

This would put the best case squeezing at about -1.25 dB, given all the above estimated losses.  

Attachment 1: scan_19-01-08_20-07-14.pdf
scan_19-01-08_20-07-14.pdf
  2293   Fri Jan 4 19:49:51 2019 awadeDailyProgressHomodyneTweek up fiber alignment and reboot of homodyne

[From before weekend]

Its not clear that in the past I had good mode overlap in the homodyne between the signal port and the local oscillator (LO).  Today I looked into that and found that some of the signal light is clipping the west PD, which would indicate that the alignment is poor.  

I used a fiber splitter to launch 1064 nm out of both the LO and signal ports from the same laser source.  Looking at the beam over lap in the near field and far field, it is pretty obvious that the beams are not coincident. They clearly hit different points on the card, not great.  

The way I configured this was to connect one leg of the 50/50 fiber splitter to the launch of the signal path and the other into the existing  2m + fiber paddle in the LO path. We should expect to see some fringing as the laser drifts or small temperature changes cause some LF drift (in this case FSR = 150 MHz). With a little tweeking of alignment, without moving any opitics I think I have them overlapping but don't see coherent fringing.  Tomorrow I'll need to remove a steering mirror.  Also problem could be caused by bad polarization, that is another thing to check.  

One way check goodness of alignment would be to drive the laser frequency with waveform much faster than thermal drift and walk the alignment of the LO + signal until the fringing peak to peak is maximized.  For the Diabolo, the PZT response is 1.95 MHz/V, 77 V pp should be enough for a full fringe of this 2 m mismatch MZ.  The only question is wether the fiber launch from the WOPO end is going to be seated exactly the same between fiber insersions and whether the slighly different MFD of the fiber will mess with the MM.

More to come.

Quote:

Today I restarted the Diabolo.  Still good on the high power generation, so that's good.

Checked the throughput of the 532 nm fiber.  Looks like from 2.04 mW at the input coliumator I get 1.05 mW at the output. This makes it 51%, ok, enough to work with.

Also checked 1064 nm alignment and MM into fiber.  There from 6.0 mW I get about 2.8 mW out from the other side.  Not the best but more than enough given that I think I only need about 2mW at most.

---

I also resurrected the homodyne. The alignment on PDs was checked.  The polarization was an arbitrary elliptical state (beam splitter is supposed to be optimal for s-pol).  I think this was left in this state because I was doing some other optimization before where I was balancing​ power using polarization.  It could also be the case that the paddle polarization controller has undergone temperature drift.  I used a beam cube to make sure that there is now only linearly polarized in the vertical direction.

With the electronic HD ballance trimmer I was able to bring the homodyne common mode rejection ratio (CMMR) down to 65 dB. This means that there is a 65 dB suppression of noise induced from RIN in the optical local oscillator.  I fine tuned this by injecting a 50 mVpp @ 3.1 kHz sine into the laser diode current modulation into (0.1A/V) on the back of the Diabolo controller and comparing the transfer function with one PD blocked and then both un-blocked.

It looks good to have a look at the impact of PZT scan on subtracted signal tomorrow.  I also need to check how much 1064 nm seed light I I can expect to get through to the WOPO copropagating with the 532 nm pump. 

 

Attachment 1: 20190104_TopViewSQZHDSetup.jpg
20190104_TopViewSQZHDSetup.jpg
  2292   Wed Jan 2 18:51:53 2019 awadeDailyProgressHomodyneTweek up fiber alignment and reboot of homodyne

Today I restarted the Diabolo.  Still good on the high power generation, so that's good.

Checked the throughput of the 532 nm fiber.  Looks like from 2.04 mW at the input coliumator I get 1.05 mW at the output. This makes it 51%, ok, enough to work with.

Also checked 1064 nm alignment and MM into fiber.  There from 6.0 mW I get about 2.8 mW out from the other side.  Not the best but more than enough given that I think I only need about 2mW at most.

---

I also resurrected the homodyne. The alignment on PDs was checked.  The polarization was an arbitrary elliptical state (beam splitter is supposed to be optimal for s-pol).  I think this was left in this state because I was doing some other optimization before where I was balancing​ power using polarization.  It could also be the case that the paddle polarization controller has undergone temperature drift.  I used a beam cube to make sure that there is now only linearly polarized in the vertical direction.

With the electronic HD ballance trimmer I was able to bring the homodyne common mode rejection ratio (CMMR) down to 65 dB. This means that there is a 65 dB suppression of noise induced from RIN in the optical local oscillator.  I fine tuned this by injecting a 50 mVpp @ 3.1 kHz sine into the laser diode current modulation into (0.1A/V) on the back of the Diabolo controller and comparing the transfer function with one PD blocked and then both un-blocked.

It looks good to have a look at the impact of PZT scan on subtracted signal tomorrow.  I also need to check how much 1064 nm seed light I I can expect to get through to the WOPO copropagating with the 532 nm pump. 

  2291   Mon Dec 24 16:46:41 2018 awadeDailyProgressWOPOMM into 532 nm patch fiber (to be hooked up to WOPO)

I had another look at MM the 532 nm light into the P3-460B-FC-2 patch cable.  After walking some lens positions and mirror pointing based on ATF:2257 I found that I could get 1.07 mW output from launch of 2.3 mW (46.5 %) through the fiber.  I expect the most I'll need is about 30 mW at the output which will put the amount of required power at about 64 mW.  There is now enough 532 nm laser light to do this and I think its within the tolerances of the fiber to have that much launched in.

One thing I did notice, on closer inspection of the fiber ends, is that there is a little bit of damage on one end (labeled A) of the P3-460B-FC-2 patch.  This is pictured below.

Yellow patch SM fiber (P3-460B-FC-2) damage at end A.

I tried a bit of gentle cleaning with some fiber cleaning cloth (Thorlabs FCC-7020)  but this appears to be a burn mark when zoomed in. Sorry doesn't capture very clear on my phone through the fiber microscope. I swapped the end that was in the fiber collimator (F240APC-532) but didn't increase the fiber launch efficiency​.  Not sure when this got damaged, but I might have exposed it to excessive power at some stage. 

I still think that the fiber through put should be sufficient, if it turns out to be a problem then re-ording shouldn't be an issue.  Thorlabs usually seems to make these fiber items next day.

Next steps

Next step is to tune up the 1064 nm alignment into the LO launch fiber and see if I can get the new PZT mounted mirror to scan as expected.  Need to figure out if there is a way to calibrate the phase on this so that I can check it is actually scanning phase (the last PZT broke and I had no idea).

Need to figure out how much 1064 nm light can be co-propagated in the patch waveguide.  The previous dismal power values might be because I made very little effort to mode match,  I'll look at what lenses I have available and see if I can boost the power through put to something respectable.

  2290   Mon Dec 24 12:07:18 2018 KojiDailyProgressWOPODiabolo: back to power with deliberate pre-misalignment

Great recovery job!

 

  2289   Mon Dec 24 10:08:49 2018 awadeDailyProgressWOPODiabolo: back to power with deliberate pre-misalignment

I think I've managed to bring the power of the Diabolo back to a usable level for WOPO.

More power with misalignment

I had another look at the alignment into the SHG.  This time I systematically trialed a range of cavity misalignments to deliberately set the pointing into the cavity to be wrong for the fast scanning mode of the control unit (cold) but to allow for the locked (hot) cavity to expand into correct alignment.  This video shows the SHG output as the cavity is locked IMG_4766.MOV,  this seems to show the expansion affecting mostly the horizontal axis. However, eventually I found that the locked cavity had a preferred misalignment on the vertical axis (in terms of the nobs that I needed to turn).  In reality its a mixture of horizontal and vertical once all the DOF are traced through the lenses and mirrors.

I was able to get a steady 160 mW of power with this new misalignment/alignment strategy.  The photos below shows the 1064 nm transmission peaks and error signal for a combination of (mostly) vertical and (a little) horizontal misalignment​ that gave greater power. 

SHG transmission peaks for alignment that delivered 160 mW
PDH error signal corresponding to above misalignment (202 mVpp amplitude)
Steady state power, looks like it peaks at much higher power
(823 mW, but couldn't find a optimal alignment and
oven temperature point that exceeded about 165 mW)

The oven operating temperature was at 110.14 C with a laser diode current of 2.102 A.  PDH gain was set to 0.56 and offset was set to 5.1.

Upping the oven temperature limit

The Diabolo SHG oven set point temperature range doesn't put the ideal phase matching temperature at the center of the available values.  In fact, the optimal power for the above-mentioned power boost was achieved pretty much of the edge of the available range (top of range is 110.26).  I popped the lid on the controller box and traced through all the PCBs to find out how the temperature range was trimmed. The front panel potentiometer (1k) is trimmed with a small set screw (blue) potentiometer located directly to the low left of the temperature knob (when viewed from the component mounting side).  I trimmed the set point temperature value so that the upper value was 111.00 C, giving an extra 0.75 C of headroom above the ideal phase matching set point.  

I found that at the very upper edge of the new range the thermal control loop started to loose some stability.  I limited the increase of temperature to the 111.00 C point and will use some caution when adjusting temperatures at these upper ranges to make sure I don't get oscillations in power.

 

 

 

 

Quote:

The strategy locking at lower laser power and optimizing phase matching temperature by slowly walking the temperature and laser power up won't work: the autolocker mode of the Diabolo SHG won't engage at lower laser power. Instead I tried a brute force approach of stepping temperature and re-locking the cavity in 0.05 C increments of set point temperature.  I've plotted the result below.  Error bars are based on standard deviation on 32 averaged measurements over 10 seconds.

I stopped at 106.00 C as  power output was not improving beyond that point. Maximum temperature is clipped at the controller at 110.26 C. Peak output of 66 mW is actually  at 110.20 C. This is counterintuitively higher than the previous set point but might be good enough to work with, as long as its stable over longer periods.

Still not clear what has changed to shift the ideal SHG over set point to 1.4 C higher than it was before.  Maybe the SHG oven needs to go even a little hotter, but we can only access tempertures up to 110.26 C. I'll take what I can get.

Some other laser settings: 

  • Injection current: 2.077 A
  • Laser crystal temperature: 26.32 C
  • SHG PDH gain: 0.5
  • PDH offset: 4.95
  • Noise eater: off

Data and plotting notebook are attached in a zip below: 20181220_SHGLockedPowerVsSetPointTemp.zip

 

  2288   Fri Dec 21 16:14:42 2018 RahulDailyProgressCryo vacuum chamberPump down results

I switched on the pumping station this morning at 10 am and roughly after 6 hours of pumping the pressure in the vacuum chamber is 8.2*10^-6 Torr. The pressure is being monitored using two gauges (wide range gauge and Ion gauge). However due to some reason the gauge controller is unable to switch on the Ion gauge (there is an inbuilt safety in the controller to protect the low pressure gauge, so this could be one of the reason).

 

I will keep the pumps running overnight and through out the weekend to monitor the pressure.

Attachment 1: Pressure.jpg
Pressure.jpg
Attachment 2: PLot1_pressure.png
PLot1_pressure.png
  2287   Fri Dec 21 15:51:27 2018 awadeDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

No I don't think so.  The temperature of the oven is important for the phase matching condition.  Its the nob you can turn so that the refractive index for both 532 nm and 1064 nm is just right so that both traveling waves stay in phase as they propagate in the non-linear crystal.  Otherwise the phase precesses in one wavelength relative to the other causing repeated amplification and de-amplification as the light propagates (rather than just amplification when they are in phase).  

Its true that temperature change induces expansion​ and dn/dt changes that will change the round trip effective length/phase. However, the PZT in the SHG should pick up the slack of from changes in the crystal round trip phase due to temperature.  We can tune the laser temperature to bring the PZT to the center of its range but at this stage it doesn't look like its railing the range of this actuator.  

My guess is that self heating affects the relative alignment of the crystal HR surface relative to the PZT mounted (curved) mirror, this small change in alignment will shift the eigen axis of the cavity and mean that the locked hot cavity will have a slightly different optimal alignment to the unlocked case.  Its much harder to walk alignment of 1064 nm while keeping cavity locked.

 

Quote:

This means that you want to make the SHG crystal longer. Is that true? If so, can you change the temperature for the optimal phase matching by tuning the 1064 crystal temperrature? I suspect you need to cool the YAG crystal, but I am not sure what is the thero-optic constant of the SHG crystal, and how much you can gain from this.

 

  2286   Fri Dec 21 15:33:26 2018 KojiDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

This means that you want to make the SHG crystal longer. Is that true? If so, can you change the temperature for the optimal phase matching by tuning the 1064 crystal temperrature? I suspect you need to cool the YAG crystal, but I am not sure what is the thero-optic constant of the SHG crystal, and how much you can gain from this.

  2285   Fri Dec 21 15:23:02 2018 RahulDailyProgressCryo vacuum chamberChamber closed for 1st pump down test

The cryo vacuum chamber was closed yesterday (Thursday) for the 1st pump down test. The figure below shows various components which was attached on to the CF flanges. Later I will post the results of the ongoing pump down test.

Attachment 1: Picture1.jpg
Picture1.jpg
Attachment 2: Picture2.jpg
Picture2.jpg
Attachment 3: Picture3.png
Picture3.png
Attachment 4: Picture4.jpg
Picture4.jpg
  2284   Thu Dec 20 19:20:50 2018 awadeDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

The strategy locking at lower laser power and optimizing phase matching temperature by slowly walking the temperature and laser power up won't work: the autolocker mode of the Diabolo SHG won't engage at lower laser power. Instead I tried a brute force approach of stepping temperature and re-locking the cavity in 0.05 C increments of set point temperature.  I've plotted the result below.  Error bars are based on standard deviation on 32 averaged measurements over 10 seconds.

I stopped at 106.00 C as  power output was not improving beyond that point. Maximum temperature is clipped at the controller at 110.26 C. Peak output of 66 mW is actually  at 110.20 C. This is counterintuitively higher than the previous set point but might be good enough to work with, as long as its stable over longer periods.

Still not clear what has changed to shift the ideal SHG over set point to 1.4 C higher than it was before.  Maybe the SHG oven needs to go even a little hotter, but we can only access tempertures up to 110.26 C. I'll take what I can get.

Some other laser settings: 

  • Injection current: 2.077 A
  • Laser crystal temperature: 26.32 C
  • SHG PDH gain: 0.5
  • PDH offset: 4.95
  • Noise eater: off

Data and plotting notebook are attached in a zip below: 20181220_SHGLockedPowerVsSetPointTemp.zip

Attachment 1: plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf
plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf
Attachment 2: 20181220_SHGLockedPowerVsSetPointTemp.zip
ELOG V3.1.3-